# Efficient Model Editing with Task Vector Bases: A Theoretical Framework and Scalable Approach

### v1

@@ B.2 Multi-task Merging Methods
Appendix C Illustration of Memory Efficiency of Task Vector Bases
In Figure 8, we illustrate the memory advantage of using task vector bases methods, from the perspective of the OOD generalization application. First, in the left figure, we observe that all storage space is used up with $T$ task vectors before clustering. In the second step, since using $m$ bases provides comparable accuracy performance to using $T$ vectors, we only need to store $m$ copies of the checkpoint. When new tasks are very different from the existing memory, as shown in Table 1, we allow $T-m$ new bases to store additional information. Thus, by reducing redundant task vectors to task vector bases, we save space for storing new vectors during domain generalization. This approach is especially beneficial when $T$ original task vectors are highly similar, and future target tasks are orthogonal to the old knowledge.
Figure 8: Workflow of using task vector bases for generation. The light gray box represents all available memory and we assume $T$ task vectors uses up the full memory as shown in the left. Colorful arrows represents different task vector bases. We use clustering and merging from left to middle figure for the bases creation step in Algorithm 1, and use the OOD generalization rule in Table 1 from middle to right which involves saving new bases into the memory. If we simply use the left figure for OOD generalization, we do not allow additional new bases which can lead to trivial performance based on Table 12.
Appendix D Details for Computer Vision Experiments
D.1 Task Vector Norm $C$
@@
D.2 Identifying Task Groups with Less Training
@@
D.3 Bases Addition
We report the full table of Table 3 in Table 8 for all possible partitions of MNIST, SVHN, EuroSAT, and RESISC45. We can draw several conclusions. First, our natural partition is almost always among the best for a 2-2 partition. In the context of multitask learning, the common wisdom is that similar tasks can be trained to leverage mutually shared information for better performance (Caruana, 1993). Additionally, by Section 4.2, at the second $\texttt{merge}_{\text{out}}$ step, it is also beneficial to combine two task vector bases with fewer conflicts or smaller $\epsilon$. These two conditions are automatically satisfied by the clustering process. Second, MS-ER is consistently better than the average of all $k=2$ and $k\in[1,4]$ partitions. By grouping, we can achieve better-than-average addition performance while using a small memory budget. Note that there is no consistent partition that always achieves the best performance for all $\texttt{merge}_{\text{in}}$ methods.
@@ D.4 Bases OOD Generalization
In Table 12, however, when we replace the target task with DTD, we do not observe significant improvement using any of the MSER partitions. This is expected because, first, based on the clustering results in Section D.2, DTD is not grouped with any of the MSER tasks, and second, in Figure 10, DTD’s cosine similarity score is consistently low.
@@
D.5 Model Merging Methods Settings
@@
For all tangent method rows in our tables, we jointly tune the corresponding datasets in one cluster, similar to MTL, and replace the standard non-linear fine-tuning with the linearized fine-tuning methods in (Ortiz-Jimenez et al., 2024). For fine-tuning-based methods in model merging, we note that they can still be used for clustering. For example, when we use Figure 11 as the affinity matrix input, the clustering result is: [(MNIST, SVHN), (SUN397, DTD, Cars), (EuroSAT), (RESISC45), (GTSRB)], where we can still observe alignment with both intermediate clustering results and the final converged clustering results in Section D.2. Additionally, the same similarity metric is used for LoRA modules in other contexts, as seen in Ostapenko et al. (2024).
@@
Appendix E Details for Language Experiments
E.1 Merging Performance on the Medium-Scale Benchmark
We include the full version of Table 6 in Section E.1. Since we employ a strong merge algorithm, Localize-and-Stitch (L&S), within our framework, we observe the advantage of the two bases L&S versions over several important baselines, including Task Arithmetic, TIES (Yadav et al., 2024), Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), AdaMerging (Yang et al., 2023), and Consensus methods (Wang et al., 2024a). Even though Mask Sharing achieves the lowest performance among all L&S variants due to its focus on memory efficiency, its multitask performance still outperforms all other non-L&S baselines.
@@ E.2 Hyperparameters
E.3 Language Tasks in the Large-Scale Benchmark

### v2

@@ B.2 Multi-task Merging Methods
Appendix C Details for Computer Vision Experiments
C.1 Task Vector Norm $C$
@@
C.2 Identifying Task Groups with Less Training
@@
C.3 Bases Addition
We report the full table of Table 4 in Table 10 for all possible partitions of MNIST, SVHN, EuroSAT, and RESISC45. We can draw several conclusions. First, our natural partition is almost always among the best for a 2-2 partition. In the context of multitask learning, the common wisdom is that similar tasks can be trained to leverage mutually shared information for better performance (Caruana, 1993). Additionally, by 4.4, at the second $\texttt{merge}_{\text{out}}$ step, it is also beneficial to combine two task vector bases with fewer conflicts or smaller $\epsilon$. These two conditions are automatically satisfied by the clustering process. Second, MS-ER is consistently better than the average of all $k=2$ and $k\in[1,4]$ partitions. By grouping, we can achieve better-than-average addition performance while using a small memory budget. Note that there is no consistent partition that always achieves the best performance for all $\texttt{merge}_{\text{in}}$ methods.
@@ C.4 Bases OOD Generalization
In Table 14, however, when we replace the target task with DTD, we do not observe significant improvement using any of the MSER partitions. This is expected because, first, based on the clustering results in Section C.2, DTD is not grouped with any of the MSER tasks, and second, in Figure 10, DTD’s cosine similarity score is consistently low.
@@
C.5 Model Merging Methods Settings
@@
For all tangent method rows in our tables, we jointly tune the corresponding datasets in one cluster, similar to MTL, and replace the standard non-linear fine-tuning with the linearized fine-tuning methods in (Ortiz-Jimenez et al., 2024). For fine-tuning-based methods in model merging, we note that they can still be used for clustering. For example, when we use Figure 11 as the affinity matrix input, the clustering result is: [(MNIST, SVHN), (SUN397, DTD, Cars), (EuroSAT), (RESISC45), (GTSRB)], where we can still observe alignment with both intermediate clustering results and the final converged clustering results in Section C.2. Additionally, the same similarity metric is used for LoRA modules in other contexts, as seen in Ostapenko et al. (2024).
@@
Appendix D Details for Language Experiments
D.1 Merging Performance on the Medium-Scale Benchmark
We include the full version of Table 8 in Section D.1. Since we employ a strong merge algorithm, Localize-and-Stitch (L&S), within our framework, we observe the advantage of the two bases L&S versions over several important baselines, including Task Arithmetic, TIES (Yadav et al., 2024), Fisher merging (Matena & Raffel, 2022), RegMean (Jin et al., 2022), AdaMerging (Yang et al., 2023), and Consensus methods (Wang et al., 2024a). Even though Mask Sharing achieves the lowest performance among all L&S variants due to its focus on memory efficiency, its multitask performance still outperforms all other non-L&S baselines.
@@ D.2 Hyperparameters
D.3 Language Tasks in the Large-Scale Benchmark
@@
D.4 Robustness of Clustering to Prompt Choice
Task Pair Similarity
(MNLI, SST2Adv) 0.0075
(MNLI, CR) 0.0080
(CR, SST2Adv) 0.0627
Table 16: Task vector similarity between SST2 trained with an adversarial 2-sentence prompt (SST2Adv), MNLI, and CR.
Due to the observation in Figure 7, one may ask if the clustering result is simply the result of prompt choice. We show that the answer is no by conducting an adversarial experiment in which we fine-tune SST-2 (Socher et al., 2013), a single-sentence binary sentiment analysis dataset, using a two-sentence task prompt template: [$S_{1}$] It was {terrible/great}. [$S_{1}$].
We present task vector similarity results in Table 16. MNLI (Williams et al., 2017) is a two-sentence natural language inference task, and CR (Hu & Liu, 2004) is a one-sentence customer review sentiment analysis task. MNLI and CR are trained with standard one- or two-sentence prompts as listed in Section D.3. We observe that even though SST-2 is trained with a two-sentence prompt, it remains much more similar to CR (another one-sentence sentiment task) than to MNLI.
This similarity pattern suggests that clustering is driven by task properties, rather than prompt structure alone. Therefore, we do not expect the number of clusters to be significantly affected by the choice of prompts.